PROBLEM #1

According to last year's report, a Filipino household spends an average of 400 pesos a day for food. Suppose you recently took random samples of 30 households, and determined how much each household spent for food each day. The results revealed a mean of 390 pesos and the standard deviation of 21.50 pesos. Using 0.01 level of significance, can it be concluded that the average amount spent per day for food of a Filipino household has decreased? Assume normality over population.


Find the critical value


1
Expert's answer
2021-06-07T01:40:10-0400

One sample t-test.

Null hypothesis "H_0:\mu=400."

Alternative hypothesis "H_a:\mu<400."

Degrees of freedom: "df=30-1=29."

Critical t-value: "t_{crit}=-2.462."

Test statistic: "t=\frac{390-400}{\frac{21.50}{\sqrt{30}}}=-2.55."

Since the test statistic lies in the critical region, reject the null hypothesis.

There is a sufficient evidence that the average amount spent per day for food of a Filipino household has decreased.